[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Jul 09, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Sun Jul 10 02:36:14 UTC 2016


HSU Shift Log for Sat Jul 09 2016 22:36:07 GMT-0400 (EDT)

Shift: R

Forecaster: R Berg

Technical Issues Summary:

ATCF issues: The scripts for the European SHIPS guidance for both Blas and  
Celia failed at 00Z. The 00Z TAFB fixes didn't import at the usual time,  
but we tried to import again around 0150 UTC and they were available.

N-AWIPS/AWIPS2 issues:

Other Techincal issues: The clocks on the walls in HSU and TAFB paused for  
about 20 minutes, but then went back to the correct time by themselves.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160710/dd841876/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list